# XY-Cut++: Advanced Layout Ordering via Hierarchical Mask Mechanism on a Novel Benchmark

### v2

@@ Abstract
Document Reading Order Recovery is a fundamental task in document image understanding, playing a pivotal role in enhancing Retrieval-Augmented Generation (RAG) and serving as a critical preprocessing step for large language models (LLMs). Existing methods often struggle with complex layouts (e.g., multi-column newspapers), high-overhead interactions between cross-modal elements (visual regions and textual semantics), and a lack of robust evaluation benchmarks. We introduce XY-Cut++, an advanced layout ordering method that integrates pre-mask processing, multi-granularity segmentation, and cross-modal matching to address these challenges. Our method significantly enhances layout ordering accuracy compared to traditional XY-Cut techniques. Specifically, XY-Cut++ achieves state-of-the-art performance (98.8 BLEU overall) while maintaining simplicity and efficiency. It outperforms existing baselines by up to 24% and demonstrates consistent accuracy across simple and complex layouts on the newly introduced DocBench-100 dataset. This advancement establishes a reliable foundation for document structure recovery, setting a new standard for layout ordering tasks and facilitating more effective RAG and LLM preprocessing.
@@ 1 Introduction
To address these challenges, we propose XY-Cut++, an advanced framework for layout ordering that incorporates three core innovations: (a) pre-mask processing to mitigate interference from high-dynamic-range elements (e.g., title, figure, table), (b) multi-granularity region splitting for the adaptive decomposition of diverse layouts, and (c) lightweight cross-modal matching leveraging minimal semantic cues. Our approach significantly outperforms traditional methods in layout ordering accuracy. Additionally, we introduce DocBench-100, a novel benchmark dataset designed for evaluating layout ordering techniques. It includes 100 pages (30 complex and 70 regular layouts) with block-level reading order annotations. Extensive experiments on DocBench-100 demonstrate that our method achieves state-of-the-art performance. Specifically, our method achieves BLEU-4 scores of 98.6(complex) and 98.9(regular), surpassing baselines by 24% on average while maintaining 1.06× FPS of geometric-only approaches.
@@ 2.1 Traditional Approaches
The XY-Cut algorithm [1] is a foundational technique in Document Reading Order Recovery that recursively divides documents into smaller regions based on horizontal and vertical projections. As illustrated in Fig. 2, while it is effective for simple layouts, it struggles with complex structures, leading to inaccuracies in reading order recovery. Specifically, the rigid threshold mechanisms of XY-Cut can introduce hierarchical errors when handling intricate layouts, resulting in suboptimal performance. To address these limitations, various enhancements have been proposed. For instance, dynamic programming optimizations [6] have been introduced to improve segmentation stability. Additionally, mask-based normalization techniques [7] have been developed to mitigate some of the challenges associated with complex layouts. However, these improvements are still insufficient for handling intricate layouts and cross-page content. In our analysis of the XY-Cut algorithm, as discussed in [6, 7] and illustrated in Fig. 3, we identified that many challenging cases arise from L-shaped inputs. A straightforward solution involves initially masking these L-shaped regions and subsequently restoring them. Upon implementing this approach, we found it to be remarkably effective while maintaining both simplicity and efficiency.
@@ 3.1 Sources and Composition
We source candidate pages from public document detection corpora (notably PP-DocLayout [19]) and MinerU [20] extraction outputs, then select pages exhibiting challenging phenomena: multi-column articles, irregular or spanning titles, interleaved figures/captions, and nested regions. The dataset comprises 100 pages split into: complex ($D_{c}$, 30 pages) and regular ($D_{r}$, 70 pages), following the prevalence in practical applications. An overview will be presented in Fig. 4.
@@ 3.3 Annotation Pipeline
We adopt a two-stage pipeline aligned with reviewer suggestions:
@@ 3.4 Statistics and Usage Protocols
4 Methodology
Our geometry–semantic fusion pipeline (Fig. 5) has four stages: (1) PP-DocLayout [19] extracts visual features and shallow semantic labels; (2) highly dynamic elements (e.g., titles, tables, figures) are pre-masked to alleviate the L-shape problem; (3) cross-layout elements are identified via global analysis, then processed with masking, real-time density estimation, and heuristic sorting; and (4) masked elements are re-mapped using nearest IoU edge-weighted margins. This geometry-aware pipeline with shallow semantics attains state-of-the-art results on DocBench-100 and OmniDocBench [3].
@@ 4.1 Pre-Mask Processing
We first identify candidate high-dynamic elements using shallow semantic labels provided by PP-DocLayout [19] and simple geometric heuristics (e.g., large area, extreme aspect ratio, or spanning multiple columns). We then construct a binary mask that temporarily excludes these elements from the subsequent multi-granularity segmentation and geometric sorting stages. During training, we randomly perturb the mask to increase robustness to imperfect detection and to teach the sorting heuristics to infer global structure without relying on a few dominant blocks. At inference time, masked elements are re-mapped back to the ordered layout using a lightweight cross-modal matching strategy: candidate masked blocks are matched to nearby unmasked regions by an IoU and margin-weighted nearest-neighbor rule that favors semantic priority (title $\succ$ text $\succ$ figure). This two-stage handle—mask first then remap—reduces L-shape artifacts while preserving final layout fidelity.
@@ 4.2 Multi-Granularity Segmentation
As shown in Fig. 7, our hybrid segmentation framework combines masking, preliminary splitting, and layout-aware refinement through three key phases:
Phase 1: Cross-Layout Detection Compute document-level median bounding box length and establish an adaptive threshold with scaling factor $\beta=1.3$:
@@
Detect cross-layout elements using dual criteria:
@@
$\mathcal{C}_{\mathrm{cross}}(B_{i})=\begin{cases}T&l_{i}>\mathcal{T}_{l}\ \land\ \sum_{j\neq i}\mathbb{I}_{\mathrm{overlap}}(B_{i},B_{j})\geq 2\\ F&\text{otherwise}\end{cases}$ (2)
Masked cross-layout elements are preserved for subsequent layout-aware splitting, while standard components undergo immediate processing.
Phase 2: Geometric Pre-Segmentation Central elements and isolated graphical components are identified through:
@@
$\mathcal{P}(B_{i})=\mathbb{I}\left(\frac{\|c_{i}-c_{\text{page}}\|_{2}}{d_{\text{page}}}\leq 0.2\right)\land(\phi_{\text{text}}(B_{i})=\infty)$ (3)
[…]
Phase 3: Density-Driven Refinement The adaptive XY-Cut algorithm dynamically selects splitting axes based on regional content density:
$\tau_{d}=\frac{\sum_{k=1}^{N_{c}}w_{k}^{(C_{c})}h_{k}^{(C_{c})}}{\sum_{k=1}^{N_{s}}w_{k}^{(C_{s})}h_{k}^{(C_{s})}}$ (4)
where $C_{c}$ denotes cross-layout elements and $C_{s}$ represents single-layout elements within region $R$. $N_{c}$ and $N_{s}$ are their respective counts.
Direction-aware splitting strategy:
$\mathcal{S}(R)=\begin{cases}\mathrm{XY\text{-}Cut}&\tau_{d}>\theta_{v}\ (\theta_{v}=0.9)\\ \mathrm{YX\text{-}Cut}&\text{otherwise}\end{cases}$ (5)
Atomic region representation:
$R_{i}=\langle x_{1}^{(i)},y_{1}^{(i)},x_{2}^{(i)},y_{2}^{(i)},C_{i}\rangle,\quad C_{i}\in\mathcal{C}_{\mathrm{type}}$ (6)
with $(x_{1}^{(i)},y_{1}^{(i)})$ and $(x_{2}^{(i)},y_{2}^{(i)})$ defining the bounding box coordinates. The content type $C_{i}$ distinguishes between cross-layout (spanning multiple grid units) and single-layout (contained within one grid unit) components.
@@ 4.3 Cross-Modal Matching
To establish cross-modal reading coherence, we propose a geometry-aware alignment framework with two core components in Fig. 7.
Multi-Stage Semantic Filtering: Orders candidates through label priority:
@@
$\begin{gathered}\mathcal{L}_{\text{order}}:\mathcal{L}_{\text{cross-layout}}\succ\mathcal{L}_{\text{title}}\succ\mathcal{L}_{\text{vision}}\succ\mathcal{L}_{\text{others}}\\ \mathcal{M}_{\text{sem}}^{(l)}=\left\{(B_{\text{p}},B_{\text{o}}^{\prime})\Bigg|\begin{aligned} &L_{\text{p}}=l\\ &\land\ L_{\text{o}}^{\prime}\succeq l\end{aligned}\right\},\quad l\in\mathcal{L}_{\text{order}}\\ \mathcal{M}_{\text{sem}}=\bigcup_{l\in\mathcal{L}_{\text{order}}}\mathcal{M}_{\text{sem}}^{(l)}\end{gathered}$ (7)
where $\mathcal{L}_{\text{vision}}$ denotes visual element detection objectives including tables, images, seals, and their corresponding titles, with $B_{p}$ representing pending bounding boxes and $B^{\prime}_{o}$ indicating ordered candidate regions.
@@
Adaptive Distance Metric: Computes joint geometric distance with early termination. Given layout element $B_{\text{p}}=(x_{1},y_{1},x_{2},y_{2})$ and candidate $B_{\text{o}}^{\prime}=(x_{1}^{\prime},y_{1}^{\prime},x_{2}^{\prime},y_{2}^{\prime})$, we define:
$D(B_{\text{p}},B_{\text{o}}^{\prime},l)=\sum_{k=1}^{4}w_{k}\cdot\phi_{k}(B_{\text{p}},B_{\text{o}}^{\prime})$ (8)
where $\phi_{k}$ encodes four geometric constraints:
  • Intersection Constraints:
@@
    $\phi_{1}=\begin{cases}1,&\text{if }\mathrm{dir}(B_{p})\neq\mathrm{dir}(B_{o}^{\prime})\lor\mathrm{IoU}_{\text{partial}}<\tau_{\text{overlap}}\\ 0,&\text{otherwise}\end{cases}$
  • Boundary Proximity:
@@
    $\phi_{2}=w_{\text{edge}}\times\begin{cases}d_{x}+d_{y},&\text{(diagonal adjacency)}\\ \min(d_{x},d_{y}),&\text{(axis-aligned)}\end{cases}$
  • Vertical Continuity:
@@
    $\phi_{3}=\begin{cases}-y_{2}^{\prime},&l\in\mathcal{L}_{\text{cross-layout}}\land y_{1}>y_{2}^{\prime}\\ y_{1}^{\prime},&\text{(baseline alignment)}\end{cases}$
  • Horizontal Ordering:
    $\phi_{4}=x_{1}^{\prime}$
We further introduce two parameterization strategies:
Dynamic Weight Adaptation. Scale-sensitive weights via dimensional analysis:
@@
$\boldsymbol{w}_{\text{edge}}=\begin{cases}[1,0.1,0.1,1]&l\in\mathcal{L}_{\text{title}}\cap\mathcal{O}_{\text{horiz}}\\ [0.2,0.1,1,1]&l\in\mathcal{L}_{\text{title}}\cap\mathcal{O}_{\text{vert}}\\ [1,1,0.1,1]&l\in\mathcal{L}_{\text{cross-layout}}\\ [1,1,1,0.1]&l\in\mathcal{L}_{\text{otherwise}}\end{cases}$ (10)
@@
Algorithm 1 The overall flow of cross-modal matching
1:Initialize $\mathcal{L}_{\text{order}}$: $\mathcal{L}_{\text{cross-layout}}\succ\mathcal{L}_{\text{title}}\succ\mathcal{L}_{\text{vision}}\succ\mathcal{L}_{\text{others}}$
2:Initialize Label Priority: $\text{doc title}\succ\text{title}\succ\text{otherwise}$
3:Initialize: $\mathcal{M}_{\text{aligned}}\leftarrow\mathcal{B}_{\text{ordered}}^{\prime}$
4:for each masked label $L\in\mathcal{L}_{\text{order}}$ do
5:  for each pending box $B\in\mathcal{B}_{\text{pending-L}}$ do
6:   Initialize $D_{\min}\leftarrow\infty$
7:   for each ordered box $B^{\prime}\in\mathcal{B}_{\text{ordered}}^{\prime}$ do
8:     $D_{\text{curr}}\leftarrow 0$
9:     for $k\leftarrow 1$ to 4 do
10:      $D_{\text{curr}}\leftarrow D_{\text{curr}}+w_{k}\cdot\phi_{k}(B,B^{\prime})$
11:      if $D_{\text{curr}}>D_{\min}$ then
12:        break $\triangleright$ filter out
13:      end if
14:     end for
15:     if $D_{\text{curr}}<D_{\min}$ then
16:      $D_{\min}\leftarrow D_{\text{curr}}$, $B^{\prime}_{\text{best}}\leftarrow B^{\prime}$
17:     end if
18:   end for
19:   $\mathcal{M}_{\text{aligned}}\leftarrow\mathcal{M}_{\text{aligned}}\cup\{(B,B^{\prime}_{\text{best}})\}$
20:  end for
21:  Reorder $\mathcal{M}_{\text{aligned}}$ by (Index, Label Priority, Y1, X1)
22:end for

### v3

@@ Abstract
Document Reading Order Recovery is a fundamental task in document image understanding, playing a pivotal role in enhancing Retrieval-Augmented Generation (RAG) and serving as a critical preprocessing step for large language models (LLMs). Existing methods often struggle with complex layouts (e.g., multi-column newspapers), high-overhead interactions between cross-modal elements (visual regions and textual semantics), and a lack of robust block-level evaluation benchmarks. We introduce XY-Cut++, an advanced layout ordering method that integrates pre-mask processing, multi-granularity segmentation, and cross-modal matching to address these challenges. Our method significantly enhances layout ordering accuracy compared to traditional XY-Cut techniques. Specifically, XY-Cut++ achieves state-of-the-art performance (98.8 BLEU overall) while maintaining simplicity and efficiency. It outperforms existing baselines by up to 24% and demonstrates consistent accuracy across simple and complex layouts on the newly introduced DocBench-100 dataset. This advancement establishes a reliable foundation for document structure recovery, setting a new standard for layout ordering tasks and facilitating more effective RAG and LLM preprocessing.
@@ 1 Introduction
[…]
@@ 2.1 Traditional Approaches
The XY-Cut algorithm [1] is a foundational technique in Document Reading Order Recovery that recursively divides documents into smaller regions based on horizontal and vertical projections. As illustrated in Figure 2, while it is effective for simple layouts, it struggles with complex structures, leading to inaccuracies in reading order recovery. Specifically, the rigid threshold mechanisms of XY-Cut can introduce hierarchical errors when handling intricate layouts, resulting in suboptimal performance. To address these limitations, various enhancements have been proposed. For instance, dynamic programming optimizations [6] have been introduced to improve segmentation stability. Additionally, mask-based normalization techniques [7] have been developed to mitigate some of the challenges associated with complex layouts. However, these improvements are still insufficient for handling intricate layouts and cross-page content. In our analysis of the XY-Cut algorithm, as discussed in [6, 7] and illustrated in Figure 3, we identified that many challenging cases arise from L-shaped inputs. A straightforward solution involves initially masking these L-shaped regions and subsequently restoring them. Upon implementing this approach, we found it to be remarkably effective while maintaining both simplicity and efficiency.
@@ 3.1 Sources and Composition
We source candidate pages from public document detection corpora (notably PP-DocLayout [19]) and MinerU [20] extraction outputs, then select pages exhibiting challenging phenomena: multi-column articles, irregular or spanning titles, interleaved figures/captions, and nested regions. The dataset comprises 100 pages split into: complex ($D_{c}$, 30 pages) and regular ($D_{r}$, 70 pages), following the prevalence in practical applications. An overview will be presented in Figure 4.
@@ 3.3 Annotation Pipeline
We adopt a two-stage pipeline:
@@ 3.4 Statistics and Usage Protocols
4 Methods
Our geometry–semantic fusion pipeline (Figure 5) has four stages: (1) PP-DocLayout [19] extracts visual features and shallow semantic labels; (2) highly dynamic elements (e.g., titles, tables, figures) are pre-masked to alleviate the L-shape problem; (3) cross-layout elements are identified via global analysis, then processed with masking, real-time density estimation, and heuristic sorting; and (4) masked elements are re-mapped using nearest IoU edge-weighted margins. This geometry-aware pipeline with shallow semantics attains state-of-the-art results on DocBench-100 and OmniDocBench [3].
@@ 4.1 Pre-Mask Processing
Our Pre-Masking mechanism is designed to reduce the interference of highly dynamic elements (e.g., figures and tables) on core block sorting. These elements can appear anywhere in documents with high positional flexibility, which may disrupt sorting logic. Specifically, we first identify such elements using shallow semantic labels from the PP-DocLayout detection model. We then build a binary mask to temporarily exclude these elements from subsequent multi-granularity segmentation and core block sorting. After multi-granularity sorting, we remap the masked elements back to the ordered layout in the cross-modal matching phase, using an IoU-weighted distance nearest-neighbor strategy. This two-stage ”mask-then-remap” pipeline effectively separates highly dynamic elements from the key sorting stage, ensuring accurate core block ordering while maintaining layout fidelity, and also reduces the previously mentioned ”L”-shaped input issues.
@@ 4.2 Multi-Granularity Segmentation
As shown in Figure 7, our hybrid segmentation framework integrates cross-layout element masking, geometric pre-segmentation, and density-driven adaptive refinement in three key phases. Each phase targets specific layout challenges, enabling robust segmentation for both regular and complex documents.
Phase 1: Cross-Layout Detection We first compute the document-level median bounding box length to establish an adaptive threshold for cross-layout elements (e.g., cross-column text spanning multiple grid units). Each content block $B_{i}$ is represented by a 4-tuple $(x_{1},y_{1},x_{2},y_{2})$, where $(x_{1},y_{1})$ is the top-left corner and $(x_{2},y_{2})$ is the bottom-right corner. Let $\{l_{i}\}_{i=1}^{N}$ denote the length of each content bounding box $B_{i}$: for horizontal-layout documents (commonly used in daily scenarios, e.g., reports, articles), $l_{i}=x_{2,i}-x_{1,i}$ (horizontal width); for vertical-layout documents (e.g., classical Chinese texts), $l_{i}=y_{2,i}-y_{1,i}$ (vertical height). $\beta=1.3$ is an empirically tuned scaling factor. The adaptive threshold $\mathcal{T}_{l}$ is:
@@
Cross-layout elements are detected via two criteria: (1) $l_{i}>\mathcal{T}_{l}$, and (2) horizontal projection overlap with at least 2 other blocks. The judgment function $\mathcal{C}_{\mathrm{cross}}(B_{i})$ is defined as:
@@
$\mathcal{C}_{\mathrm{cross}}(B_{i})=\begin{cases}1&l_{i}>\mathcal{T}_{l}\ \land\ \sum_{j\neq i}\mathbb{I}_{\mathrm{overlap}}(B_{i},B_{j})\geq 2\\ 0&\text{otherwise}\end{cases}$ (2)
Here, $\mathbb{I}_{\mathrm{overlap}}(B_{i},B_{j})$ is an indicator function (1 for horizontal projection overlap, 0 otherwise). Detected cross-layout elements are masked for subsequent layout-aware splitting, while single-layout elements (within one grid unit) are processed immediately to avoid interference.
Phase 2: Geometric Pre-Segmentation This phase aims to identify central content elements, including body titles (e.g., chapter and section titles), and isolated graphical components (e.g., figures and tables). These elements are leveraged to perform coarse-grained partitioning of the original page into multiple non-overlapping sub-regions, denoted as $R$. Each $R$ is subsequently subjected to density-driven adaptive refinement in Phase 3 to achieve fine-grained segmentation. Such elements are classified based on geometric features, formalized by the judgment function $\mathcal{P}(B_{i})$:
@@
$\mathcal{P}(B_{i})=\mathbb{I}\left(\frac{\|c_{i}-c_{\text{page}}\|_{2}}{d_{\text{page}}}\leq 0.2\right)\land(\phi_{\text{text}}(B_{i})=\infty),$ (3)
Where:
$c_{i}=(cx_{i},cy_{i})$: Center of $B_{i}$, computed as $cx_{i}=\frac{x_{1,i}+x_{2,i}}{2}$, $cy_{i}=\frac{y_{1,i}+y_{2,i}}{2}$;
$c_{\text{page}}$: Page center coordinate;
$\|c_{i}-c_{\text{page}}\|_{2}$: Euclidean distance between $B_{i}$ and page center;
$d_{\text{page}}$: The normalization factor $d_{\text{page}}$ is determined by the block type:
$d_{\text{page}}=\begin{cases}W_{\text{page}},&\text{if }r_{i}<3,\\ H_{\text{page}},&\text{if }r_{i}\geq 3,\end{cases}\quad\text{where}\quad r_{i}=\frac{w_{i}}{h_{i}}=\frac{x_{2,i}-x_{1,i}}{y_{2,i}-y_{1,i}},$
and $W_{\text{page}}$ , $H_{\text{page}}$ are the page width and height, respectively.
$\mathbb{I}(\cdot)$: Indicator function (1 if condition holds, 0 otherwise);
$\phi_{\text{text}}(B_{i})$: Minimum Euclidean distance from $B_{i}$ to the nearest text block; $\phi_{\text{text}}(B_{i})=\infty$ means $B_{i}$ is isolated (no adjacent text);
Target elements $B_{i}$: Body titles (chapter/section titles), visual components (e.g., Figure, Table), all with high layout flexibility.
Elements with $\mathcal{P}(B_{i})=1$ are marked as isolated and added to the mask set. Their coordinates are used to split the page into non-overlapping sub-regions $R$, laying the groundwork for recursive refinement.
Phase 3: Density-Driven Refinement To achieve fine-grained segmentation, we apply the adaptive XY-Cut algorithm to each coarse-grained sub-region $R$ from Phase 2. This algorithm dynamically selects the splitting axis (horizontal/vertical) based on regional content density of the current sub-region $R$. The density metric $\tau_{d}$ quantifies the ratio of cross-layout element area to single-layout element area within $R$, calculated as:
$\tau_{d}=\frac{\sum_{k=1}^{N_{c}}w_{k}^{(C_{c})}h_{k}^{(C_{c})}}{\sum_{k=1}^{N_{s}}w_{k}^{(C_{s})}h_{k}^{(C_{s})}},$ (4)
where $C_{c}$ and $C_{s}$ are cross-layout and single-layout element sets in $R$; $N_{c}$, $N_{s}$ are their counts. $w_{k}^{(C_{c})}h_{k}^{(C_{c})}$ and $w_{k}^{(C_{s})}h_{k}^{(C_{s})}$ represent the area of the $k$-th element in each set, respectively.
A higher $\tau_{d}$ indicates a denser distribution of cross-layout elements (typically horizontal content). We set a density threshold $\theta_{v}=0.9$ to determine the splitting direction, as defined by the strategy function $\mathcal{S}(R)$:
$\mathcal{S}(R)=\begin{cases}\mathrm{XY\text{-}Cut}&\tau_{d}>\theta_{v}\\ \mathrm{YX\text{-}Cut}&\text{otherwise}\end{cases}$ (5)
Here, XY-Cut refers to prioritizing horizontal splitting (along the Y-axis): we compute the Y-axis projection histogram of region $R$, find the position $y_{\text{cut}}$ with the minimum projection value (i.e., the ”gap” between content blocks), and split $R$ into upper and lower sub-regions. YX-Cut prioritizes vertical splitting (along the X-axis) using the same logic on the X-axis projection histogram.
The recursive splitting process terminates when each sub-region contains only one non-masked content block. Each final atomic region $R_{i}$ is represented as a 7-tuple, encapsulating spatial coordinates, content type, semantic label, and positional index:
$R_{i}=\langle x_{1}^{(i)},y_{1}^{(i)},x_{2}^{(i)},y_{2}^{(i)},C_{i},\text{Label}_{i},\text{Index}_{i}\rangle$ (6)
where $(x_{1}^{(i)},y_{1}^{(i)})$ and $(x_{2}^{(i)},y_{2}^{(i)})$ denote the top-left and bottom-right coordinates of $R_{i}$; $\mathcal{C}_{\mathrm{type}}=\{\text{cross-layout},\text{single-layout}\}$, with $C_{i}\in\mathcal{C}_{\mathrm{type}}$ indicating the content type of $R_{i}$; $\text{Label}_{i}$ represents the semantic label (e.g., title, figure, paragraph) of $R_{i}$; and $\text{Index}_{i}$ denotes the positional index of $R_{i}$ in the document layout. This 7-tuple representation provides comprehensive information to support subsequent cross-modal matching.
@@ 4.3 Cross-Modal Matching
To establish reading coherence consistent with human habits, we propose a geometry-aware cross-modal alignment framework, consisting of multi-stage semantic filtering and adaptive distance metric modules (Figure 7). This framework fuses semantic priority and geometric constraints to realize accurate ordering of atomic regions.
Multi-Stage Semantic Filtering: This module realizes the restoration of masked elements and optimizes the ordering sequence by leveraging label priority, ensuring high-semantic elements (e.g., cross-layout text, titles) are prioritized in the matching and restoration process. We first define core sets and a global label priority sequence, then formalize the matching logic to sequentially restore masked elements into the ordered sequence.
The multi-stage semantic filtering module restores masked elements by leveraging a global label priority sequence $\mathcal{L}_{\text{order}}$. The process operates on three core sets: the pre-ordered atomic region set $S$, the masked element set $\mathcal{M}$, and the dynamically updated target sequence $T$ (initialized as $T=S$). The mathematical formulation is:
@@
$\begin{gathered}\mathcal{L}{{}_{\text{order}}}:\mathcal{L}{{}_{\text{cross-layout}}}\succ\mathcal{L}{{}_{\text{title}}}\succ\mathcal{L}{{}_{\text{vision}}}\succ\mathcal{L}{{}_{\text{others}}}\\ \mathcal{F}(B{{}_{\text{p}}},T,l_{\text{current}})=\begin{cases}1,&\exists B_{\text{o}}\in T,\ \mathcal{L}{{}_{\text{order}}}(l(B{{}_{\text{o}}}))\succ\mathcal{L}{{}_{\text{order}}}(l_{\text{current}})\\ 0,&\text{otherwise}\end{cases}\\ \mathcal{M}^{(l_{\text{current}})}{{}_{\text{sorted}}}=\{(B{{}_{\text{p}}},B{{}_{\text{best}}})\mid B{{}_{\text{p}}}\in\mathcal{M},\ l(B{{}_{\text{p}}})=l_{\text{current}},\ \mathcal{F}=1,\ B{{}_{\text{best}}}\in T\}\\ T=T\cup\mathcal{M}^{(l_{\text{current}})}{{}_{\text{sorted}}},\quad\forall l_{\text{current}}\in\mathcal{L}{{}_{\text{order}}}\end{gathered}$ (7)
Here, the binary matching function $\mathcal{F}$ verifies for a pending element $B_{\text{p}}$ (with label $l_{\text{current}}$) whether a higher-priority candidate exists in $T$. For each $B_{\text{p}}$ that satisfies $\mathcal{F}=1$, the specific optimal anchor $B_{\text{best}}\in T$ is determined by minimizing a joint geometric distance (detailed in the following Adaptive Distance Metric). The set $\mathcal{M}^{(l_{\text{current}})}_{\text{sorted}}$ then contains all such matched pairs $(B_{\text{p}},B_{\text{best}})$ for the current label. These pairs are subsequently integrated into $T$, and the process iterates through $\mathcal{L}_{\text{order}}$, ensuring high-priority elements are restored first.
This multi-stage semantic filtering strategy effectively eliminates irrational matching pairs (e.g., visual elements paired with footnotes) and reinforces the semantic coherence of the final ordered sequence, laying a solid foundation for downstream cross-modal alignment.
@@
Adaptive Distance Metric: We design a joint geometric distance metric with early termination to find the optimal matching position for each $B_{\text{p}}$ in $\mathcal{M}_{\text{sem}}$. Given a pending layout element $B_{\text{p}}=(x_{1},y_{1},x_{2},y_{2})$ and an ordered candidate $B_{\text{o}}=(x_{1}^{\prime},y_{1}^{\prime},x_{2}^{\prime},y_{2}^{\prime})$, the distance $D(B_{\text{p}},B_{\text{o}},l)$ (parameterized by $B_{\text{p}}$’s label $l$) is a weighted sum of four geometric constraints $\phi_{k}$:
$D(B_{\text{p}},B_{\text{o}},l)=\sum_{k=1}^{4}w_{k}\cdot\phi_{k}(B_{\text{p}},B_{\text{o}}),$ (8)
where $w_{k}$ are adaptive weights (detailed later), and $\phi_{1}\sim\phi_{4}$ encode different geometric constraints to capture spatial relationships:
  • Intersection Constraints ($\phi_{1}$): Filters invalid pairs by layout direction and overlap. The overlap threshold is set to $\tau_{\text{overlap}}=0.3$.
@@
    $\phi_{1}=\begin{cases}1,&\text{if }\mathrm{direction}(B_{p})\neq\mathrm{direction}(B_{o}^{\prime})\lor\mathrm{IoU}_{\text{projection}}<\tau_{\text{overlap}}\\ 0,&\text{otherwise}\end{cases}$ (9)
  • Boundary Proximity ($\phi_{2}$): Measures spatial adjacency. Weighted by $w_{\text{edge}}$, it uses center distances, prioritizing axis-aligned adjacency.
@@
    $\phi_{2}=w_{\text{edge}}\times\begin{cases}d_{x}+d_{y},&\text{(diagonal adjacency)}\\ \min(d_{x},d_{y}),&\text{(axis-aligned)}\end{cases}$ (10)
  • Vertical Continuity ($\phi_{3}$): Optimizes vertical ordering.
@@
    $\phi_{3}=\begin{cases}-y_{2}^{\prime},&l\in\mathcal{L}_{\text{cross-layout}}\land y_{1}>y_{2}^{\prime}\\ y_{1}^{\prime},&\text{(baseline alignment)}\end{cases}$ (11)
  • Horizontal Ordering ($\phi_{4}$): Follows left-to-right reading logic via $B_{\text{o}}$ left boundary.
    $\phi_{4}=x_{1}^{\prime}$ (12)
We introduce two parameterization strategies to optimize the metric’s adaptability:
Dynamic Weight Adaptation. To ensure the four geometric constraint distances do not interfere with each other, we design scale-sensitive weights with staggered scaling levels. The weight vector $\boldsymbol{w}_{k}=[w_{1},w_{2},w_{3},w_{4}]$ is formulated as:
@@
$\boldsymbol{w}_{\text{edge}}=\begin{cases}[1,0.1,0.1,1]&l\in\mathcal{L}_{\text{title}}\cap\mathcal{O}_{\text{horizontal}}\\ [0.2,0.1,1,1]&l\in\mathcal{L}_{\text{title}}\cap\mathcal{O}_{\text{vertical}}\\ [1,1,0.1,1]&l\in\mathcal{L}_{\text{cross-layout}}\\ [1,1,1,0.1]&l\in\mathcal{L}_{\text{otherwise}}\end{cases}$ (14)
where $\mathcal{O}_{\text{horizontal}}$ and $\mathcal{O}_{\text{vertical}}$ denote the horizontal layout set and vertical layout set, respectively; $l$ is the semantic label of the element. This weight design assigns differentiated weights based on label characteristics and layout orientations, further weakening mutual interference between constraints and improving matching stability across various scenarios.
@@
Algorithm 1 Geometry-Aware Cross-Modal Matching with Semantic Filtering
1:Sorted anchor set $\mathcal{S}$, masked element set $\mathcal{M}$, distance weights $w_{1},w_{2},w_{3},w_{4}$
2:Matching result set $\mathcal{T}$
3:Initialize global label priority: $\mathcal{L}_{\text{order}}:\mathcal{L}_{\text{cross-layout}}\succ\mathcal{L}_{\text{title}}\succ\mathcal{L}_{\text{vision}}\succ\mathcal{L}_{\text{other}}$
[…]
5:for each semantic label $l$ in $\mathcal{L}_{\text{order}}$ (descending priority) do
6:  for each pending box $B_{p}\in\mathcal{M}$ with label $l$ do $\triangleright$ Check $\mathcal{F}(B_{p},\mathcal{T},l)$
7:   $D_{\min}\leftarrow\infty$, $B_{\text{best}}\leftarrow\text{null}$
8:   for each ordered box $B_{o}$ in $\mathcal{T}$ do $\triangleright$ Search candidates in $\mathcal{T}$ for $\mathcal{F}$
9:     $D_{\text{curr}}\leftarrow 0$
10:     for $k\leftarrow 1$ to 44 do
11:      $D_{\text{curr}}\leftarrow D_{\text{curr}}+w_{k}\cdot\phi_{k}(B_{p},B_{o})$
12:      if $D_{\text{curr}}>D_{\min}$ then
13:        break $\triangleright$ Early termination
14:      end if
15:     end for
16:     if $D_{\text{curr}}<D_{\min}$ then
17:      $D_{\min}\leftarrow D_{\text{curr}}$, $B_{\text{best}}\leftarrow B_{o}$
18:     end if
19:   end for
20:   if $B_{\text{best}}\neq\text{null}$ then $\triangleright$ Valid match found for $B_{p}$
21:     $\mathcal{T}\leftarrow\mathcal{T}\cup\{(B_{p},B_{\text{best}})\}$ $\triangleright$ $\mathcal{M}_{\text{sorted}}^{(l)}\leftarrow\mathcal{M}_{\text{sorted}}^{(l)}\cup\{(B_{p},B_{\text{best}})\}$
22:   end if
23:  end for
24:  Sort $\mathcal{T}$ by: Label priority of $B_{p}$ (desc) $\succ$ Index of $B_{\text{best}}$ (asc) $\succ$ $y_{1}$ of $B_{p}$ (asc) $\succ$ $x_{1}$ of $B_{p}$ (asc)
25:end for
26:return $\mathcal{T}$